# Simulating discrete-time quantum walk with urn model

### v1

@@ Abstract
and assign phases $\eta_{0,0}(0)=\phi_{0}$, $\eta_{0,1}(0)=\phi_{1}$ so as to encode the initial coin–position state of Eq.27.
@@
Figure 2: Snapshot of the trajectories of the marked ball in the urn model ($N=10^{9}$) of the 1D Hadamard quantum walk, initialized in the state $\frac{1}{\sqrt{2}}\lvert 0\rangle-\frac{i}{\sqrt{2}}\lvert 1\rangle$, shown at time $t=200$
@@
At every site $x$ with $N_{x,0}(t)+N_{x,1}(t)>0$, we apply a $\mathbb{T}_{c}$ process on the two urns labeled by $(x,0)$ and $(x,1)$ with $\tilde{N}$ and $\eta_{i}^{\prime}$ as given in Eq. (24) and Eq. (26) respectively. This is a stochastic mixing rule of the two urns at each site that mirrors the quantum “coin” operation. Concretely, urn‐occupancies and phases are updated to new values $N^{\prime}_{x,c}(t)$ and $\eta^{\prime}_{x,c}(t)$ so as to reproduce, in the large–$N$ limit, the SU(2) coin operation of Eq. (4).
@@
After the $\mathbb{T}_{c}$, all balls in urn $c=0$ at site $x$ move to urn 00 of site $x+1$, while those in $c=1$ move to urn $1$ of site $x-1$. Phases are also carried along unchanged:
@@
and all other urn‐occupancies reset to zero.
@@
After $t$ steps of iterating $\mathbb{T}_{c}$ and $\mathbb{S}$, we perform the measurement process $\mathbb{M}$: we identify the urn $(x_{\rm m},c_{\rm m})$ containing the marked ball, move all balls into that urn, and record the site index $x_{\rm m}$. The resulting histogram of recorded $x_{\rm m}$ over many trials reproduces the probability distribution of the quantum walk at time $t$ in the large $N$ limit [Fig. 1].
@@
The dynamics can be viewed from a different perspective also: by iterating $\mathbb{T}_{c}$ and $\mathbb{S}$ for $t$ steps, the marked ball undergoes a classical random walk [Fig. 2] whose large–$N$ distribution converges to that of the quantum walker’s position distribution $|\psi(x,t)|^{2}$.
@@
In this work, we have developed an urn model to simulate the dynamics of 1D discrete-time quantum walks with a generalized coin. The urn model can be systematically generalized to higher-dimensional graphs. In the two-dimensional case, for example, the walker requires a $4\times 4$ coin operation acting on a two-qubit state, representing four possible directions: up, down, left, and right. Consequently, the urn model must be extended to include a transformation operation $\mathbb{T}$ that acts between four distinct urns, with each site in the two-dimensional lattice hosting four urns. By carefully extending these ideas, the urn model can be generalized to arbitrary higher-dimensional graphs or even complex networks, where each node may be associated with multiple urns corresponding to the available directions or connections. So, this approach offers a versatile and intuitive framework for simulating a broad range of quantum dynamical systems. Our future work will focus on formalizing this methodology and exploring its application to increasingly complex topologies.
The strength of the present model lies in its simplicity and the valuable insights it offers. Successful extensions of this approach could enable the design of DTQW-based algorithms through purely urn-based reasoning and may open new pathways for the efficient quantum simulation of complex classical systems.
Quantum walks have found successful applications across a wide range of domains, including quantum search algorithms [23], community detection in complex networks [24], modeling energy transfer in photosynthesis [25, 26, 27], modeling quantum mechanics in curved space-time [28, 29], to name a few. The split-step quantum walk is particularly notable for its ability to simulate relativistic quantum dynamics [30, 31]. Discrete-time quantum walks have also proven to be powerful tools for simulating topological insulators [32, 33], enabling the study of topologically protected edge states and phase transitions through precise control over coin and shift parameters. Non-Hermitian [34] and open quantum walks extends the standard quantum walk framework to incorporate loss, gain, and environmental interactions, making them powerful tools for modeling quantum active particles [35] and non-equilibrium dynamics in open quantum systems. The present study establishes a connection between the urn model and all these domains where quantum walks are effectively applied.
Properties such as the first-passage problem, stochastic resetting and time of arrival- traditionally studied in the context of classical stochastic processes, can also be explored in quantum walks [36, 37, 38] and other discrete-time quantum processes [39]. Such attempts have been made in the literature using the tools of quantum mechanics. We believe the present work will offer a more accessible framework and open new avenues for future exploration in this direction using the language and tools of the urn model and the corresponding random walks of the marked ball.
@@
The author thanks Nicolas Gisin, Parongama Sen, and Arghya Das for valuable discussions.
Appendix:
Suppose In the $\mathbb{P}$ process, we set the parameters to match the quantum model: $\eta_{i}=\phi_{i}$ and $\rho_{0}=r_{0}^{2}$. Consequently, the urn occupation number $N_{0}$ is chosen randomly as either
$N_{0}=\lfloor r_{0}^{2}N\rfloor\quad\text{or}\quad N_{0}=\lceil r_{0}^{2}N\rceil$
with equal probability $1/2$. Given that the total number of balls is $N$, the complement is
$N_{1}=N-N_{0}.$
We can always write:
$r_{0}^{2}N=\lfloor r_{0}^{2}N\rfloor+\epsilon=\lceil r_{0}^{2}N\rceil-\epsilon^{\prime},$
where $10\leq\epsilon,\epsilon^{\prime}<$ 1. Moreover, if $r_{0}^{2}N\notin\mathbb{Z}$, then $\epsilon+\epsilon^{\prime}=1$. This implies:
$N_{0}=r_{0}^{2}N+O(1),\quad N_{1}=r_{1}^{2}N+O(1).$
Multiplying gives:
$N_{0}N_{1}=(r_{0}^{2}N+O(1))(r_{1}^{2}N+O(1))=r_{0}^{2}r_{1}^{2}N^{2}+O(N).$
We write $O(f(N))$ to indicate any quantity bounded in absolute value by $Cf(N)$ for some constant $C>0$.
Let us define:
$A:=r_{0}^{2}r_{1}^{2}N^{2},\quad B:=O(N),$
so that:
$\sqrt{N_{0}N_{1}}=\sqrt{A+B}.$
Now factor out the leading-order part:
$\sqrt{A+B}=\sqrt{A}\cdot\sqrt{1+\frac{B}{A}}.$
Since $B=O(N)$ and $A=O(N^{2})$, we have:
$\frac{B}{A}=O\left(\frac{1}{N}\right).$
Using a Taylor expansion:
$\sqrt{1+\frac{B}{A}}=1+\frac{1}{2}\cdot\frac{B}{A}+O\left(\frac{1}{N^{2}}\right),$
which gives:
$\sqrt{A+B}=\sqrt{A}\left(1+O\left(\frac{1}{N}\right)\right)=r_{0}r_{1}N+O(1).$
From the Hadamard-like transformation, we know:
${r_{0}^{\prime}}^{2}=\frac{1}{2}\left(r_{0}^{2}+r_{1}^{2}+2r_{0}r_{1}\cos(\phi_{0}-\phi_{1})\right),$
so multiplying by $N$:
${r_{0}^{\prime}}^{2}N=\frac{1}{2}\left(r_{0}^{2}N+r_{1}^{2}N+2r_{0}r_{1}N\cos(\phi_{0}-\phi_{1})\right).$
Substituting $N_{0}=r_{0}^{2}N+O(1)$, $N_{1}=r_{1}^{2}N+O(1)$, and $\sqrt{N_{0}N_{1}}=r_{0}r_{1}N+O(1)$, we obtain:
${r_{0}^{\prime}}^{2}N=\frac{1}{2}\left(N_{0}+N_{1}+2\sqrt{N_{0}N_{1}}\cos(\phi_{0}-\phi_{1})\right)+O(1).$ (31)
If we define:
$\tilde{N}_{0}:=\frac{1}{2}\left(N_{0}+N_{1}+2\sqrt{N_{0}N_{1}}\cos(\phi_{0}-\phi_{1})\right),$
then:
${r_{0}^{\prime}}^{2}N=\tilde{N}_{0}+O(1).$
Hence, one can equivalently write:
${r_{0}^{\prime}}^{2}N=\lfloor\tilde{N_{0}}\rfloor+O(1)$ (32)
${r_{0}^{\prime}}^{2}N=\lceil\tilde{N_{0}}\rceil+O(1)$ (33)
${r_{0}^{\prime}}^{2}N=\frac{\lfloor\tilde{N_{0}}\rfloor+\lceil\tilde{N_{0}}\rceil}{2}+O(1)$ (34)
Thus, in the asymptotic limit:
$\lim_{N\to\infty}\frac{\lfloor\tilde{N}_{0}\rfloor+\lceil\tilde{N}_{0}\rceil}{2N}={r_{0}^{\prime}}^{2}.$ (35)
Using a similar approach, one can show that in the limit $N\to\infty$, $\eta_{i}^{\prime}(N_{0},N_{1},\phi_{0},\phi_{1})$ of Eq.21 approaches to $\phi_{i}^{\prime}$. Also one can prove using $\tilde{N_{0}}$ and $\eta_{i}^{\prime}$ of Eq. 24 and 26 respectively that the probability of obtaining 0 and 1 in the urn model $\mathbb{M}$ process approaches corresponding probabilities in Quantum version for generalised $SU(2)$ coin operation.

### v3

@@ Abstract
and assign phases $\eta_{0,0}(0)=\phi_{0}$, $\eta_{0,1}(0)=\phi_{1}$ so as to encode the initial coin–position state of Eq.29. All other phase tags are set to zero, i.e.,
$\eta_{x,c}=0\quad\text{for}\quad(x,c)\neq(0,0),(0,1).$
@@
Figure 4: Snapshot of the trajectories of the marked ball in the boxes-and-balls model ($N=10^{9}$) of the 1D Hadamard quantum walk, initialized in the state $\frac{1}{\sqrt{2}}\lvert 0\rangle-\frac{i}{\sqrt{2}}\lvert 1\rangle$, shown at time $t=200$
@@
At every site $x$ with $N_{x,0}(t)+N_{x,1}(t)>0$, we apply a $\mathbb{T}_{c}$ process on the two boxes labeled by $(x,0)$ and $(x,1)$ with $\tilde{N}_{x,0}$, $\eta_{x,k}$ and $\eta_{x,k}^{\prime}$ will play the role of $\tilde{N}_{0}$, $\phi_{k}$ and $\eta_{k}^{\prime}$ respectively as given in Eq. (35) and Eq. (36). This is a stochastic mixing rule of the two boxes at each site that mirrors the quantum “coin” operation. Concretely, box‐occupancies and phases are updated to new values $N^{\prime}_{x,c}(t)$ and $\eta^{\prime}_{x,c}(t)$ so as to reproduce, in the large–$N$ limit, the SU(2) coin operation of Eq. (4).
@@
After the $\mathbb{T}_{c}$, all balls in box $c=0$ at site $x$ move to box 0 of site $x+1$, while those in $c=1$ move to box $1$ of site x1x-1. Phases are also carried along unchanged:
@@
and all other box‐occupancies reset to zero.
@@
After $t$ steps of iterating $\mathbb{T}_{c}$ and $\mathbb{S}$, we perform the measurement process $\mathbb{M}$: we identify the box $(x_{\rm m},c_{\rm m})$ containing the marked ball, move all balls into that box, and record the site index $x_{\rm m}$. The resulting histogram of recorded $x_{\rm m}$ over many trials reproduces the probability distribution of the quantum walk at time $t$ in the large $N$ limit [Fig. 3].
@@
The dynamics can be viewed from a different perspective also: by iterating $\mathbb{T}_{c}$ and $\mathbb{S}$ for $t$ steps, the marked particle undergoes a classical random walk [Fig. 4] whose large–$N$ distribution converges to that of the quantum walker’s position distribution $|\psi(x,t)|^{2}$.
Active spin models [21, 22, 23, 24, 25] describe collections of self-propelled particles carrying an internal two-state degree of freedom (spin), whose dynamics combines stochastic spin flips with biased motion on a lattice. Such models are known to exhibit rich nonequilibrium phenomena, including collective motion, phase separation, and long-range order in low dimensions. The primary motivation for such models is twofold: first, to identify the minimal ingredients required for collective motion; and second, to understand how nonequilibrium driving circumvents equilibrium constraints on ordering and phase transitions. Building on the box–ball formulation of the discrete-time quantum walk, we introduce an active spin model obtained directly from this framework.
In this construction, each ball is reinterpreted as an active particle and is endowed with an Ising spin variable $s=\pm 1$ and a real-valued phase $\eta_{x}^{\,s}$, inherited from the box index $c\in\{0,1\}$ and the associated phase tag $\eta_{x,c}$ of the box–ball representation. The dynamics consists of two coupled processes: (i) spin-flip and phase update events that modify the internal state as well as the phase of a particle, and (ii) active hopping on the lattice, in which the direction and rate of motion depend on the spin. The spin-flip and hopping rates are chosen so that the particle density produces the probability distribution of the discrete time quantum walker when the number of particles are sufficiently large.
The explicit definition of the spin variables, flip rates, and hopping rules is introduced below. For the moment, the essential point is that the classical active dynamics is engineered to mirror the unitary coin–shift structure of the quantum walk.
We consider $N$ particles carrying Ising spins $s=\pm 1$ on a one-dimensional lattice of $L$ sites. In addition to the spin degree of freedom, each particle carries a real-valued phase $\eta_{x}^{\,s}$ that depends only on its lattice position $x$ and spin $s$. Each particle hops to its right and left neighboring sites at rates $D(1+\varepsilon s)$ and $D(1-\varepsilon s)$, respectively. There is no exclusion between particles, and we denote by $n_{x}^{\pm 1}$ the number of $\pm$ spins on site $x$, so that the local density is given by $\rho_{x}=n_{x}^{+1}+n_{x}^{-1}$. The particles also flip their spins: on site $x$ each spin flips at rate
$R_{x,s}=\frac{1}{2}(1+\frac{s\delta_{x}}{\lvert\delta_{x}\rvert})\frac{\delta_{x}}{n_{x}^{s}},$
Here $\delta_{x}=\tilde{n}_{x}^{+1}-n_{x}^{+1}$, where $\tilde{n}_{x}^{+1}$ is defined by the same expression as $\tilde{N}_{0}$ in Eq. (35), with the local variables $n_{x}^{+1}$, $n_{x}^{-1}$, $\eta_{x}^{+1}$, and $\eta_{x}^{-1}$ playing the roles of $N_{0}$, $N_{1}$, $\phi_{0}$, and $\phi_{1}$, respectively.
[…]
The updated phases $\eta_{x}^{+1}$$^{\prime}$ and $\eta_{x}^{-1}$$^{\prime}$ correspond to $\eta_{x,0}^{\prime}$ and $\eta_{x,1}^{\prime}$ in the box-and-balls construction.
${\eta_{x}^{s}}^{\prime}=\mathrm{atan2}\left(\frac{\mathbb{B}_{s}}{\mathbb{A}_{s}}\right),$ (36)
where,
$\mathbb{B}_{s}=\sqrt{\frac{n_{x}^{+1}}{N}}\cos\theta\sin(\xi+\eta_{x}^{+1})+s\sqrt{\frac{n_{x}^{-1}}{N}}\sin\theta\sin(\zeta+\eta_{x}^{-1}),$
$\mathbb{A}_{s}=\sqrt{\frac{n_{x}^{+1}}{N}}\cos\theta\cos(\xi+\eta_{x}^{+1})+s\sqrt{\frac{n_{x}^{-1}}{N}}\sin\theta\cos(\zeta+\eta_{x}^{-1}).$
For $D=0.5$ and $\varepsilon=1$, i.e., when the hopping is completely biased, if the initial particle configuration matches the corresponding initial configuration of the box-and-ball model, the resulting density distribution reproduces the probability distribution of the discrete-time quantum walk in the large-$N$ limit. This behavior has been confirmed by direct numerical simulations.
Recently, it has been shown in the literature that suitably designed active matter systems can exhibit emergent features characteristic of quantum dynamics, including interference-like effects [26], the ability to tunnel through potential barriers [27], Bose–Einstein condensation [28], Schrödinger-type dynamics in polar liquids [29], spin–orbit coupling [30], topological effects [31] and Hall viscosities [32, 33] . Theses allows for the simulation of quantum phenomena in classical, controlled environments. Motivated by these developments, the present construction aims to provide a minimal, lattice-based active model whose collective dynamics reproduces the behavior of a discrete-time quantum walk. Since quantum walks are known to be universal simulators of quantum dynamics (any quantum circuit can be encoded in an appropriate quantum-walk architecture), this mapping highlights the expressive power of active matter as a classical platform for simulating quantum-like phenomena.
@@
In this letter, we have developed classical interacting particle models to simulate the dynamics of 1D discrete-time quantum walks with a generalized coin. The classical models can be systematically generalized to higher-dimensional graphs. In the two-dimensional case, for example, the walker requires a $4\times 4$ coin operation acting on a two-qubit state, representing four possible directions: up, down, left, and right. Consequently, the box-ball model must be extended to include a transformation operation $\mathbb{T}$ that acts between four distinct boxes, with each site in the two-dimensional lattice hosting four boxes. By carefully extending these ideas, the box model can be generalized to arbitrary higher-dimensional graphs or even complex networks, where each node may be associated with multiple boxes corresponding to the available directions or connections. So, this approach offers a versatile and intuitive framework for simulating a broad range of quantum dynamical systems. Our future work will focus on formalizing this methodology and exploring its application to increasingly complex topologies.
The strength of the present model lies in its simplicity and the valuable insights it offers. Successful extensions of this approach could enable the design of DTQW-based algorithms through purely particle-base reasoning. Various active spin models can themselves simulate a wide variety of collective dynamics. As this generalised active spin model can reflect the behavior of quantum walk, they can lead to the development of novel quantum algorithms for simulating various collective phenomena in active systems, as well as quantum-inspired classical strategies. in this regard it is important to notice that the present work demonstrates that Discrete-time quantum walks are themselves a quantum algorithm for simulating a class of classical active spin dynamics in the limit where the number of particles tends to infinity.
Quantum walks have found successful applications across a wide range of domains [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46], from quantum search algorithms [34] to simulating relativistic quantum dynamics [41, 42] and modeling quantum active particles [46]. The present study establishes a connection between the generalised active spin model and all these domains where quantum walks are effectively applied.
In contrast to classical random walks, quantum walks do not permit a complete record of the walker’s trajectory due to the measurement-induced collapse of the quantum state, which disrupts its coherent evolution. As a result, traditional notions like first-passage time (FPT) are not directly applicable, and the first-detected-passage time (FDPT) becomes the relevant quantity of interest [47, 48]. However, within the box model framework of DTQWs, it becomes possible to define both FPT and FDPT meaningfully. This dual capability offers a novel platform for examining first passage times in quantum system.
@@
Acknowledgments: The author thanks Nicolas Gisin, Parongama Sen, and Arghya Das for valuable discussions.